. Given the Production Function Q = 72X + 15X2 - X3, where Q =Output and X=Input (5

marks)

a) What is the Marginal Product (MP) when X = 8?

b) What is the Average Product (AP) when X = 6?

c) At what value of X will Q be at its maximum?

d) At what value of X will Diminishing Returns set in?


1
Expert's answer
2021-05-24T12:51:47-0400

Production function:

"Q=72x+15x^2-x^3"


(a) Marginal Product "(MP)=\dfrac{dQ}{dx}=72+30x-3x^2"


   at x=8,

      "MP=72+30(8)-3(8)^2=72+240-192=120"


(b) Average Product "(AP)=\dfrac{Q}{X}=72+15X-X^2"

    

    At x=6,

      "AP=72+15(6)-(6)^2=72+90-36=126"


(c) Maximizing Q:

    "\dfrac{dQ}{dx}=0\Rightarrow 72+30x-3x^2=0\\[9pt]x^2-10x-24=0\\(x-12)(x+2)=0"


 x=12,-2


At x=12, Output maximized.


(d)For Diminishing returns:


    "\dfrac{d MP}{dx}=\dfrac{d^2Q}{dx^2}=30-6X<0"


       "\dfrac{30}{6}<X\\[9pt] 5<X"


  After 5 level of inputs diminishing returns set in.